The firm is expected to grow forever at a rate of 2 percent per year. The current rate of interest is 5 percent, which is expected to remain unchanged for the foreseeable future. What is the value of the firm if it is currently earning profits of $1,000?

  • $26,500

  • $35,000

  • $31,000

  • $30,000

  • None of the options.
Answer #1

Current profits = $ 1000

Interest rate, k = 5%

growth rate , g = 2%

Value of the firm = Profit/(k-g) = 1000/(5% - 2%) = $ 33,333.33

So Answer is None of the options